OpenStudy (anonymous):

find the slope of the lines passing through the points (-3,4) and (2,-1)

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (math_man21):

got the answer

hartnn (hartnn):

denominator = 2-(-3) = ... ?

hartnn (hartnn):

\(\Large \dfrac{-1-4}{2-(-3)} = \dfrac{-5}{5} = ...\)

OpenStudy (anonymous):

-1

hartnn (hartnn):

\(\huge \checkmark \)